For the following reactions:
Ca(s) + 2HCl(aq) → CaCl2(aq) + H2(g) ΔHrxn = -543.0 kJ/mol
CaS(s) + 2HCl(aq) → CaCl2(aq) + H2S(g) ΔHrxn = -81.1 kJ/mol
CaS(s) → Ca(s) + 1/8 S8(s) ΔHrxn = +482.4 kJ/mol
What is the enthalpy of the reaction
H2(g) + 1/8 S8(s) → H2S(g) ?
1
Expert's answer
2020-05-01T14:41:00-0400

"\Delta" Hreaction = "\sum" "\nu" * Δ Hproducts - ν * Δ Hstarting materials

ΔH (Ca) = ΔH (H2) = 0


For firs reaction find ΔH(CaCl2):

ΔH(CaCl2) - 2ΔH(HCl) = -543 kJ/mol

ΔH(CaCl2) = 2ΔH(HCl) - 543 kJ/mol


For second reaction find ΔH(H2S):


ΔH(H2S) + ΔH(CaCl2) - ΔH(CaS) - 2ΔH(HCl) = -81.1 kJ/mol


Insert a value of ΔH(CaCl2):


ΔH(H2S) + 2ΔH(HCl) - 543 kJ/mol - ΔH(CaS) - 2ΔH(HCl) = -81.1 kJ/mol

ΔH(H2S) - 543 kJ/mol - ΔH(CaS) = -81.1 kJ/mol

ΔH(H2S) = ΔH(CaS) + 461.9 kJ/mol


For third reaction find ΔH(S8):


1/8ΔH(S8) - ΔH(CaS) = 482.4 kJ/mol

1/8ΔH(S8) = ΔH(CaS) + 482.4 kJ/mol


Enthalpy of the reaction H2(g) + 1/8 S8(s) → H2S(g) is:


Δ Hreaction = ΔH(H2S) - 1/8ΔH(S8) = ΔH(CaS) + 461.9 kJ/mol - ΔH(CaS) - 482.4 kJ/mol = -20.5 kJ/mol